OpenStudy (inkyvoyd):

2.2 3, 5, 7 Boyce seperable diff eq

OpenStudy (inkyvoyd):

3. y'+y^2 sin x=0

OpenStudy (inkyvoyd):

y^-2 y'=-sin x \(\int y^{-2}dy=\int-\sin x dx\) \(-y^{-1}=\cos x+C\) \(y=1/(-\cos x+C)\)

OpenStudy (inkyvoyd):

5. y' = cos^2 x cos^2 (2y)

OpenStudy (irishboy123):

\(3 y'+y^2 sin x=0\) \(\color{red}{3}y^{-2} y'=-sin x\)